# A Generalized Sylvester-Fermat-Torricelli problem with application in disaster relief operations by UAVs

### v1

@@ Abstract
Unmanned Aerial Vehicles (UAVs) have emerged as valuable tools in such scenarios. In particular, a fleet of UAVs can be dispatched from a mobile station to the affected area to facilitate data collection and establish temporary communication networks. Nevertheless, real-world deployment of UAVs faces several challenges, with adverse weather conditions—especially wind—being among the most significant. To address this, we develop a novel mathematical framework to determine the optimal location of a mobile UAV station while explicitly accounting for the heterogeneity of the UAVs and the effect of wind. Our approach extends a classical single-facility location problem by incorporating heterogeneous dynamic sets that represent varying UAV speed. In particular, we generalize the Sylvester problem to introduce the Sylvester-Fermat-Torricelli (SFT) problem, which captures complex factors such as wind influence, UAV heterogeneity, and back-and-forth motion within a unified framework. The proposed framework enhances the practicality of UAV-based disaster response planning by accounting for real-world factors such as wind and UAV heterogeneity. Experimental results demonstrate that it can reduce wasted operational time by up to 84%, making post-disaster missions significantly more efficient and effective.
@@
Generalized Sylvester problem, Unmanned aerial vehicle (UAV), Generalized Sylvester-Fermat-Torricelli problem, Nonsmooth optimization, Convex analysis.
@@ 1.1 Introduction and Motivation
Natural and human-made disasters worldwide can have devastating consequences, leading to significant loss of life and extensive destruction, often resulting in substantial economic and social costs. For instance, in 2023, the Emergency Events Database (EM-DAT) recorded 399 disasters worldwide, leading to 86,473 fatalities and affecting approximately 93.1 million people [10]. Given the significant costs associated with disasters, disaster management has become a critical field focused on minimizing these costs. Effective disaster management strategies can be categorized into pre-disaster activities (prevention and preparedness) and post-disaster (response and recovery) ones [8].
@@
In post-disaster actions, one of the most critical tasks is data collection. This data includes the locations of victims trapped under debris, the extent of road destruction and accessibility, and the damage levels across different parts of the region of interest. In this context, UAVs offer significant advantages, including rapid area surveying, ease of deployment, low-cost maintenance, and the ability to reach inaccessible locations [18], making them highly effective for data gathering. Moreover, since natural disasters can damage or render network infrastructure inoperable [8], UAVs can also serve as a vital tool for establishing temporary communication networks. Therefore, several studies have explored the use of UAVs to provide wireless or Delay-Tolerant Networks (DTNs) in such scenarios [40, 20, 1, 8, 14, 9, 42, 43].
@@
While UAVs can be valuable for providing network coverage, their main drawback is limited energy capacity, requiring them to return periodically for recharging [40]. Also, unlike other vehicles, UAVs are susceptible to wind due to their lighter weight, smaller size, lower flight altitude, and lower speed [11]. These limitations create significant challenges in deploying UAVs for post-disaster network provisioning. [8] proposed a mobile first-response UAV station with a fleet of heterogenous UAVs that has a long-distance communication antenna, an electricity generator and a system for automatic UAV battery recharging. Motivated by this idea, this paper aims to develop a reliable mathematical solution for determining the optimal location of a mobile UAV station, considering UAV heterogeneity and the impact of wind, which is common in real-world scenarios.
@@
This problem can be regarded as a single-facility location problem, previously studied in the context of drone-assisted delivery, where drones launch from a truck, deliver parcels, and subsequently return to the truck. There are studies [26, 3, 38, 2] in the literature that examines the location of launch points (location of truck) for drones in a delivery network where a truck and multiple drones operate in tandem. [39] considered a tandem delivery strategy like [26] to use in the healthcare application. However, To the best of our knowledge, little attention has been given to addressing the impact of wind on drone performance. This factor cannot be ignored as it causes delays in delivery time and increases energy consumption. We also tried to consider the effect of wind in other applications that are search and rescue [13], and package delivery [17].
@@
[6] provides a comprehensive review of drone-assisted delivery and highlights future research directions. They emphasize that weather conditions, particularly wind, should be accurately accounted for in future studies. Moreover, as noted in [7], existing drone-based research has not fully leveraged advancements from non-drone-related studies, including the use of heterogeneous fleets. These observations underscore a broader challenge: modern facility location problems in UAV applications demand frameworks that can simultaneously address environmental uncertainties and fleet heterogeneity. To bridge this gap, our method explicitly considers a heterogeneous fleet of UAVs and the effect of wind, making it more aligned with real-world applications. This forms another aspect of the motivation for this work in UAV applications.
1.2 Contribution
@@
To address the challenges discussed in the previous section, namely the effects of wind and UAV heterogeneity, we focus on single-facility location problems. These problems have strong theoretical foundations and have been studied in the field of convex analysis, particularly in the context of the Generalized Fermat-Torricelli and Generalized Sylvester problems. To provide a clear foundation for our approach, we first present a brief overview of these classical problems. In the 19th century, English mathematician Sylvester (1814–1897) introduced the smallest enclosing circle problem, which seeks to determine the smallest circle capable of enclosing a given set of points in a plane [41]. This problem is formulated as a minmax location problem and generalized in several ways in the literature. The French mathematician Fermat (1601–1665) proposed an optimization problem that seeks to identify a point in the plane that minimizes the total distance to three specified points. This problem was solved by the Italian mathematician and physicist Torricelli (1608–1647) and is known as the Fermat-Torricelli problem. These problems continue to attract attention both for their mathematical elegance and for their significant practical applications, particularly in location science.
The literature contains several studies on these problems.Variants of the Generalized Sylvester problem have been explored in prior research [23, 31, 29, 30, 34], while the Generalized Fermat-Torricelli problems have been investigated in [27, 24, 28, 21, 19, 15, 33, 12]. Based on the existing literature, we summarize the contributions and generalization aspects of the present work; to the best of our knowledge, these contributions span both theoretical advancements and practical applications, including:
@@
  • Introducing the generalized Sylvester-Fermat-Torricelli (SFT) problem, a novel formulation not previously explored in the literature. The generalized SFT problem is a min–max optimization problem that incorporates multiple Minkowski gauge functions rather than a single one. This generalization provides sufficient flexibility to account for UAV heterogeneity (e.g., varying speeds) and the influence of wind as a source of weather uncertainty. Furthermore, the model extends from individual points to convex sets, which is particularly useful when UAVs are required to reach a target area rather than a specific point. Finally, in this new min–max framework, the maximum of functions—each potentially representing the sum of several Minkowski gauge functions—is minimized. This approach enables modeling of UAV round-trip movements, a feature essential for realistic operational scenarios.
@@
  • Presenting an extended version of the generalized SFT problem, as well as the generalized Fermat–Torricelli and generalized Sylvester problems, incorporating two types of Minkowski gauge functions: the set-based Minkowski gauge function (2.6) and the maximal Minkowski gauge function (3.6). This extension enables modeling of scenarios in which UAVs enter a target area from the point closest to a facility and exit from the point farthest from it. In Section 6, the practical applicability of this extended formulation in real-world operations is demonstrated.
@@
  • Developing a single facility location model for disaster relief using UAVs that explicitly accounts for wind effects and UAV fleet heterogeneity, thereby enhancing its applicability to real-world scenarios.
2 Essential Definitions, Problem Formulation, and Convex Analysis Tools
@@
In this section, we begin by presenting key concepts and tools from convex analysis, which serve as a foundational framework for addressing the challenges posed by the nonsmooth nature of our model introduced in (2.7). These tools will be instrumental in formulating and analyzing the variational properties of the problem, as well as in developing effective solution approaches. Next, we present the essential definitions and the formal problem formulation. We also emphasize the key generalizations introduced in our model, which extend the capabilities beyond those of existing generalized Sylvester problems in the literature. These extensions allow us to address more complex scenarios that are closer to real-world applications.
@@ 2.1 Tools of Convex Analysis
In this section, we present key definitions and results from convex analysis that are essential for the subsequent sections.
A subset $\Omega$ of $\mathbb{R}^{q}$ is convex if $\lambda x+(1-\lambda)y\in\Omega$ for all $x,y\in\Omega$ and $\lambda\in(0,1)$. Let $f:\Omega\to\bar{\mathbb{R}}:=(-\infty,\infty]$ be an extended-real-valued function defined on a convex set $\Omega$. Then $f$ is convex on $\Omega$ if
@@
Let $\Omega\subset\mathbb{R}^{q}$ be a nonempty convex set. For $\bar{x}\in\Omega$, the Normal Cone to $\Omega$ at $\bar{x}$ is
@@
If $\bar{x}\notin\Omega$, we have $N(\bar{x};\Omega)=\emptyset$ by convention.
The following proposition demonstrates that convexity is preserved under certain important operations (the proof and details are provided in Proposition 1.44 of [22]).
@@
Let $f_{i}:\mathbb{R}^{q}\to\overline{\mathbb{R}}$ be convex functions for all $i=1,\ldots,m$. Then the following functions are also convex:
@@
The following lemma shows that the maximum of strictly convex functions is strictly convex.
Lemma 2.2.
Let $f_{i}:\mathbb{R}^{q}\to\bar{\mathbb{R}}$ be strictly convex for all $i=1,2,\dots,m$. Then, the function $f_{m}=\max_{i}f_{i}$ is strictly convex.
@@
We begin by proving the result for the case $m=2$, i.e., $f_{m}(x)=\max\{f_{1}(x),f_{2}(x)\}$. For $\in\mathbb{R}^{q}, xyx\neq y$, and $\lambda\in(0,1)$, there are two cases:
Case 1: Suppose $f_{m}(\lambda x+(1-\lambda)y)=f_{1}(\lambda x+(1-\lambda)y)$. Since $f_{1}$ is strictly convex, we have:
$f_{1}(\lambda x+(1-\lambda)y)<\lambda f_{1}(x)+(1-\lambda)f_{1}(y)\leq\lambda f_{m}(x)+(1-\lambda)f_{m}(y),$
so
$f_{m}(\lambda x+(1-\lambda)y)<\lambda f_{m}(x)+(1-\lambda)f_{m}(y).$
Case 2: Similarly, if $f_{m}(\lambda x+(1-\lambda)y)=f_{2}(\lambda x+(1-\lambda)y)$, then the same inequality holds using the strict convexity of $f_{2}$.
Therefore, $f_{m}$ is strictly convex for $m=2$. The general result for $m>2$ follows by induction.
One of the central concepts in convex analysis is the notion of a subgradient, which extends the idea of a derivative to nonsmooth convex functions. Let $f:\mathbb{R}^{q}\to\bar{\mathbb{R}}$ be a convex function, a vector $v\in\mathbb{R}^{q}$ is called a subgradient of $f$ at $\bar{x}$, $\bar{x}\in\operatorname{dom}f$, if the following inequality holds:
@@
The set of all subgradients of $f$ at $\bar{x}$ is known as the subdifferential of $f$ at $\bar{x}$, and is denoted by $\partial f(\bar{x})$. For completeness, we present two useful rules for computing subdifferential.
Theorem 2.3 (max rule, see Theorem 3.27 in [22]).
Let $f_{i}:\mathbb{R}^{q}\to\overline{\mathbb{R}}$, $i=1,\ldots,m$, be convex functions with $\bar{x}\in\bigcap_{i=1}^{m}\textnormal{dom}\ f_{i}$, and let
$f(x)=\ \max\ \{f_{i}(x)\ |\ i=1,\dots,m\}.$
Then, assuming that $f_{i}$ are continuous at $\bar{x}$ for $i=1,\dots,m$, we have
$\partial f(\bar{x})=\ \textnormal{co }\{\bigcup_{i\in I(\bar{x})}\partial f_{i}(\bar{x})\ \}$
where $I(\bar{x})=\{i=1,\dots,m\ |\ f(\bar{x})=f_{i}(\bar{x})\}$ (the “active” functions at $\bar{x}$).
The theorem says that the subgradient of $f$ at $\bar{x}$ is the convex hull of the subgradients of all functions that achieve the maximum at $\bar{x}$. In the formal statement, dom denotes the domain of a function, i.e., the set of all valid input values for which the function is defined; co denotes the convex hull of a set. Specifically, for a set $\Omega\subset\mathbb{R}^{q}$,
$\operatorname{co}\Omega:=\left\{\sum_{i=1}^{k}\lambda_{i}x_{i}\,\middle|\,x_{i}\in\Omega,\ \lambda_{i}\geq 0,\ \sum_{i=1}^{k}\lambda_{i}=1,\ k\in\mathbb{N}\right\}.$ (2.4)
Theorem 2.4 (The subdifferential of the sum equals the sum of the subdifferential, derived from Theorem 3.18 in [22] by induction.).
Let $f_{i}:\mathbb{R}^{q}\rightarrow\mathbb$$\bar${R}} for $i=1,2,\dots,m$ be convex functions. The following equality holds
$\partial(\sum_{i=1}^{m}f_{i})(\bar{x})=\sum_{i=1}^{m}\partial f_{i}(\bar{x})\ \ \ \textnormal{for every}\ \ \ \bar{x}\in\bigcap_{i=1}^{m}\textnormal{dom}f_{i}$
under the qualification condition:
$\bigcap_{i=1}^{m}\textnormal{ri(dom}f_{i})\neq\emptyset.$
This rule allows us to break down a complicated function into simpler parts in optimization. In the formal statement, ri denotes the relative interior of a set, as defined in (8.2), a subset of points in the set that are in the interior relative to the affine hull of (or space spanned by) the set.

### v2

# New Location Science Models with Applications to UAV-Based Disaster Relief

@@ Abstract
Unmanned Aerial Vehicles (UAVs) have emerged as valuable tools in such scenarios. In particular, a fleet of UAVs can be dispatched from a mobile station to the affected area to facilitate data collection and establish temporary communication networks. Nevertheless, real-world deployment of UAVs faces several challenges, with adverse weather conditions–especially wind–being among the most significant. To address such challenges, we develop a novel mathematical framework to determine the optimal location of a mobile UAV station while explicitly accounting for the heterogeneity of the UAVs and the effect of wind. Our approach extends well-known single-facility location models by incorporating heterogeneous dynamic sets that represent different UAV operating speeds. In particular, we extend the classical Fermat-Torricelli and Sylvester problems to introduce the generalized Sylvester-Fermat-Torricelli (SFT) model and its remarkable specifications that capture complex factors such as wind influence, UAV heterogeneity, and back-and-forth motion within a unified framework. The proposed framework enhances the practicality of UAV-based disaster response planning by accounting for major real-world factors including wind and UAV heterogeneity. Experimental results demonstrate that our developments can reduce wasted operational time by up to 84% while making post-disaster missions significantly more efficient and effective in practical settings.
@@
location science, Facility location problems, Unmanned aerial vehicle (UAV). Generalized Sylvester-Fermat-Torricelli models, Convex analysis, Nonsmooth optimization
@@ 1.1 State-of-the-Art and Motivations
Natural and human-made disasters worldwide can have devastating consequences leading to significant loss of life and extensive destruction, often resulting in substantial economic and social costs. For instance, in 2023, the Emergency Events Database (EM-DAT) recorded 399 disasters worldwide causing 86,473 fatalities and affecting approximately 93.1 million people [CRED2024]. Given the significant costs associated with disasters, disaster management has become a critical field focused on minimizing these costs. Effective disaster management strategies can be categorized into pre-disaster activities (prevention and preparedness) and post-disaster (response and recovery) ones [erdelj2017wireless].
@@
In post-disaster actions, one of the most critical tasks is data collection. This data includes the locations of victims trapped under debris, the extent of road destruction and accessibility, and the damage levels across different parts of the region of interest. In this context, UAVs offer significant advantages, including rapid area surveying, ease of deployment, low-cost maintenance, and the ability to reach inaccessible locations [lyu2023unmanned], making them highly effective for data gathering. Moreover, since natural disasters can damage or render network infrastructure inoperable [erdelj2017wireless], UAVs can also serve as a vital tool for establishing temporary communication networks. Therefore, several studies have explored the use of UAVs to provide wireless or Delay-Tolerant Networks (DTNs) in such scenarios [shakhatreh2019uavs, matracia2021topological, 8490824, erdelj2017wireless, khan2022emerging, erdelj2017help, wang2021disaster, 8641424].
@@
While UAVs can be valuable for providing network coverage, their main drawback is limited energy capacity requiring them to return periodically for recharging [shakhatreh2019uavs]. Also, unlike other vehicles, UAVs are susceptible to wind due to their lighter weight, smaller size, lower flight altitude, and lower speed [Gianfelice2022]. These limitations create significant challenges in deploying UAVs for post-disaster network provisioning. [erdelj2017wireless] proposed a mobile first-response UAV station with a fleet of heterogeneous UAVs that has a long-distance communication antenna, an electricity generator, and a system for automatic UAV battery recharging. Motivated by this idea, we aim here to develop a reliable mathematical model for determining the optimal location of a mobile UAV station, considering UAV heterogeneity and the impact of wind, which are common factors in real-world scenarios.
@@
Such a model can be regarded as a single-facility location problem previously studied in the context of drone-assisted delivery, where drones launch from a truck, deliver parcels, and subsequently return to the truck. There are studies [mourelo2016optimization, chang2018optimal, salama2020joint, betti2023dispatching] in the literature that examine the location of launch points (location of truck) for drones in a delivery network, where a truck and multiple drones operate in tandem. [scott2017drone] considered a tandem delivery strategy like [mourelo2016optimization] to use in the healthcare application. However, to the best of our knowledge, little attention has been given to addressing the impact of wind on drone performance. This factor cannot be ignored as it causes delays in delivery time and increases energy consumption. We also tried to consider the effect of wind in other applications that are search and rescue [kazemdehbashi2025algorithm], and package delivery [liu2023routing].
@@
The paper by [dukkanci2024facility] provides a comprehensive review of drone-assisted delivery and highlights future research directions. It emphasizes that weather conditions–particularly wind–should be accurately accounted for in future studies. Moreover, as noted in [dukkanci2024facility_EJOR], existing drone-based research has not fully leveraged advancements from non-drone-related studies including the use of heterogeneous fleets. These observations underscore a broader challenge: modern facility location problems in UAV applications demand frameworks that can simultaneously address environmental uncertainties and fleet heterogeneity. To bridge this gap, our method explicitly considers a heterogeneous fleet of UAVs and the effect of wind while making it more aligned with real-world applications. This forms yet another aspect of our motivation for this work.
1.2 Contributions
@@
To address the challenges discussed in the previous section, namely the effects of wind and UAV heterogeneity, we focus on single-facility models of location science. Problems of this type have strong theoretical foundations and have been studied using various techniques. To provide a clear foundation for our approach, we first present a brief overview of the classical problems in this area. The French mathematician Pierre de Fermat (1601–1665) proposed an optimization problem that seeks to identify a point in the plane that minimizes the total distance to three other plane points. This problem was solved by the Italian physicist Augusto Torricelli (1608–1647) and is now known as the Fermat-Torricelli problem. In the 19th century, the English mathematician James Joseph Sylvester (1814–1897) introduced the smallest enclosing circle problem, which seeks to determine the smallest circle capable of enclosing a given set of points in a plane [sylvester1857question]. This problem is formulated as a minimax location problem and has been extended in several ways in the literature. Over the centuries, these classical problems have continued to attract attention of researchers and practitioners for their mathematical elegance and significant practical applications. Various generalizations of Fermat-Torricelli, Sylvester and related problems (Steinitz, Weber, etc.) have been introduced and studied in location science and practical areas by using different approaches and methods; see, e.g., [borwein1990, boltyanski1999, kuhn1976, plastria2009asymmetric, wesolowsky1993, weiszfeld37, martini2002fermat], and the references therein.
More recent years have witnessed the development of novel approaches to location science models based on variational ideas with the usage of generalized differential tools of convex and variational analysis. To the best of our knowledge, such an approach was first suggested in [mordukhovich2011applications] for the newly introduced generalized Fermat-Torricelli problem and then was largely developed in [reich2022, kupitz2013fermat, mordukhovich2019fermat, mordukhovich2013smallest, nam2014constructions, nam2013generalized, jahn2015minsum, nam2014nonsmooth, nam2017minimizing, mordukhovich2023easy] and other publications for various models of location science with algorithmic applications.
The major contributions of our work, including theoretical advances and practical applications, are as follows.
@@
  • We introduce and investigate the generalized Sylvester-Fermat-Torricelli (SFT) problem, a novel model of location science not previously explored in the literature. The generalized SFT problem belongs to the area of minimax optimization that incorporates multiple Minkowski gauge functions rather than a single one. This generalization provides sufficient flexibility to account for UAV heterogeneity and the influence of wind as a source of weather uncertainty. Furthermore, the model extends from individual target points to convex sets, which is particularly useful when UAVs are required to reach a target area rather than a specific point. Finally, in this new minimax framework, the maximum of functions–each potentially representing the sum of several Minkowski gauge functions–is minimized. This approach enables modeling of UAV round-trip movements, a feature that is highly essential for realistic operational scenarios.
@@
  • We further introduce and discuss an extended version of the generalized SFT model, as well as extended counterparts of the generalized Fermat–Torricelli and Sylvester problems, by incorporating two types of Minkowski gauge functions: the set-based Minkowski gauge (2.5) and the maximal set-based Minkowski gauge (3.11). These extensions enable us modeling scenarios in which UAVs enter a target area from the point closest to a facility and exit from the point farthest from it. In Section 6, the practical applicability of the extended formulations in real-world operations is demonstrated with providing extensive numerical experiments.
@@
  • We develop a single facility location model for disaster relief by using UAVs that explicitly accounts for wind effects and UAV fleet heterogeneity, which therefore enhance its applicability to real-world scenarios.
The reminder of the paper is organized as follows. In Section 2, we first review some background material from convex analysis needed below and then formulate the novel generalize Sylvester-Fermat-Torricelli model with its major special cases. Section 3 is devoted to the introduction and a detailed study of the set-based Minkowski gauge function, which plays a crucial role in the paper. Section 4 verifies well-posedness issues (existence and uniqueness of optimal solutions) for the generalized SFT. In Section 5, we discuss some extended and weight versions of the generalized SFT. Section 6 contains the major applications of our theoretical developments to disaster relief operations by using UAVs with presenting the results of numerical experiments. In Section 7, we summarize the main achievements of the paper and discuss some topics of our future research. The Appendix (Section 8) presents some additional material of its own interest, which is used in the proofs of the obtained results.
2 Basic Definitions and Problem Formulation
@@
This section is started with reviewing key concepts and tools of convex analysis, which serve as a foundational framework for addressing the major challenges posed by the nonsmooth nature of our model introduced in (2.6). These tools will be instrumental in revealing and analyzing variational properties of the model, as well as in developing effective solution approaches. Then essential definitions and the problem formulation are presented and discussed. We also emphasize the key features introduced in our model, which extend the capabilities beyond those of existing models in location science. These extensions allow us to address more complex scenarios that are closer to real-world applications.
@@ 2.1 Tools of Convex Analysis
In this subsection, we present key definitions and results of convex analysis that are essential for the main results of the paper; see, e.g., the books [rockafellar2015convex, mordukhovich2022] for more details, additional material, and references. Throughout the paper, we use the standard notation from the aforementioned books. Recall that the symbol $\Omega_{1}\subset\Omega_{2}$ means that the set $\Omega_{1}$ is smaller than or equal to the set $\Omega_{2}$.
A nonempty subset $\Omega$ of $\mathbb{R}^{q}$ is convex if $\lambda x+(1-\lambda)y\in\Omega$ for all $x,y\in\Omega$ and $\lambda\in(0,1)$. An extended-real-valued function $f:\Omega\to\bar{\mathbb{R}}:=(-\infty,\infty]$ is convex on a convex set $\Omega$ if
@@
Given a nonempty convex set $\Omega\subset\mathbb{R}^{q}$ and a point $\bar{x}\in\Omega$, the normal cone to $\Omega$ at $\bar{x}$ is
@@
If $\bar{x}\notin\Omega$, we have $N(\bar{x};\Omega):=\emptyset$ by convention.
The convexity of functions is preserved under important operations.
@@
Let $f,f_{i}:\mathbb{R}^{q}\to\overline{\mathbb{R}}$ be convex functions for all $i=1,\ldots,m$. Then the following functions are also convex:
@@
The next result shows that the maximum of strictly convex functions is strictly convex.
Proposition 2.2.
Let $f_{i}:\mathbb{R}^{q}\to\overline{\mathbb{R}}$ be strictly convex function for all $i=1,\dots,m$. Then the maximum function $f(x):=\max_{1\leq i\leq m}f_{i}(x),\;x\in\mathbb{R}^{q}$, is strictly convex on $\mathbb{R}^{q}$.
@@
Consider first the case where $m=2$, i.e., $f(x)=\max\{f_{1}(x),f_{2})(x)\}$. Fix $\in\mathbb{R}^{q}, xyx\neq y$, and $\lambda\in(0,1)$ and suppose that $f(\lambda x+(1-\lambda)y)=f_{1}(\lambda x+(1-\lambda)y)$. It follows from the strict convexity of $f_{1}$ that
$f_{1}(\lambda x+(1-\lambda)y)<\lambda f_{1}(x)+(1-\lambda)f_{1}(y)\leq\lambda f(x)+(1-\lambda)f(y),$
which tells us therefore that
$f(\lambda x+(1-\lambda)y)<\lambda f(x)+(1-\lambda)f(y).$
The remaining case where $f(\lambda x+(1-\lambda)y)=f_{2}(\lambda x+(1-\lambda)y)$ is treated similarly. For an arbitrary $m\in\mathbb{N}$, we proceed by induction. ∎
One of the central concepts in convex analysis is the subdifferential of convex functions, which extends the classical derivative/gradient notion to nondifferentiable (nonsmooth) functions. Given an extended-real-valued convex function $f:\mathbb{R}^{q}\to\overline{\mathbb{R}}$ finite at $\bar{x}$ (i.e., with $\bar{x}\in{\rm dom}\,f$), a subgradient $v\in\mathbb{R}^{q}$ of $f$ at $\bar{x}$ is defined by
@@
The collection of all subgradients $v$ of $f$ at $\bar{x}$ is called the subdifferential of $f$ at this point and is denoted by $\partial f(\bar{x})$. It is well known that the subdifferential of any convex function is nonempty at every point where the function is continuous.
The subdifferential of convex functions enjoys comprehensive calculus with respect to operations that keep the convexity. For the reader’s convenience, we present the two fundamental calculus results used in what follows. The first result is the subdifferential maximum rule.
Theorem 2.3.
Let $f_{i}:\mathbb{R}^{q}\to\overline{\mathbb{R}}$, $i=1,\ldots,m$, be convex functions with $\bar{x}\in\bigcap_{i=1}^{m}\textnormal{dom}\ f_{i}$. Assume that all $f_{i}$ are continuous at $\bar{x}$ and consider the maximum function
$f(x)=\ \max\big\{f_{i}(x)\;\big|\;i=1,\dots,m\big\},\quad x\in\mathbb{R}^{q}.$
Then the subdifferential of $f$ at $\bar{x}$ is calculated by
$\partial f(\bar{x})=\ \textnormal{co }\Big\{\bigcup_{i\in I(\bar{x})}\partial f_{i}(\bar{x})\Big\},$
where $I(\bar{x}):=\{i=1,\dots,m\ |\ f(\bar{x})=f_{i}(\bar{x})\}$ is the collection of active indexes at $\bar{x}$, and where the symbol “co” signifies the convex hull if the set in question.
Theorem 2.3 tells us that the subdifferential of $f$ at $\bar{x}$ is the convex hull of the subgradients of the functions that achieve the maximum at $\bar{x}$. Recall that the convex hull of a set $\Omega\subset\mathbb{R}^{n}$ is
$\operatorname{co}\Omega:=\left\{\sum_{i=1}^{k}\lambda_{i}x_{i}\,\middle|\,x_{i}\in\Omega,\ \lambda_{i}\geq 0,\ \sum_{i=1}^{k}\lambda_{i}=1,\ k\in\mathbb{N}\right\},$
where $k$ can be bounded by $n+1$ due to the classical Carathéodory theorem.
The following calculus result used below is the fundamental subdifferential sum rule known also as the Moreau-Rockafellar theorem.
Theorem 2.4.
Let $f_{i}:\mathbb{R}^{q}\rightarrow\mathbb$$\overline${R}} for $i=1,\dots,m$ be convex functions such that all but one of $f_{i}$ are continuous at $\bar{x}\in\bigcap_{i=1}^{m}\textnormal{dom}\ f_{i}$. Then we have
$\partial\Big(\sum_{i=1}^{m}f_{i}\Big)(\bar{x})=\sum_{i=1}^{m}\partial f_{i}(\bar{x}).$